OpenStudy (anonymous):

tan 2(teta) - cot 2(teta)=0

OpenStudy (anonymous):

Is that \[\tan ^{2}\theta - \cot ^{2}\theta = 0 \] ?

OpenStudy (anonymous):

tan 2θ - cot 2θ =0

OpenStudy (abb0t):

\[\tan(\theta) = \frac{ \sin(\theta ) }{ \cos(\theta) }\]\[\cot(\theta) = \frac{ \cos(\theta) }{ \sin(\theta) }\]

OpenStudy (anonymous):

then cross multiplication?

OpenStudy (anonymous):

Ok, try \[\tan(2\theta) = \cot(2\theta) --> \frac{ \sin(2\theta) }{ \cos(2\theta) } = \frac{ \cos(2\theta) }{ \sin(2\theta) } \] Cross multiply to obtain \[\sin ^{2} (2\theta) = \cos ^{2}(2\theta)\] I used various techniques from here. The method is up to you.

OpenStudy (anonymous):

then bring cos to produce tan??

OpenStudy (anonymous):

i got it...:)

OpenStudy (anonymous):

give me the final answer please

OpenStudy (anonymous):

\[\sin ^{2}(2\theta) = \cos ^{2}(2\theta) -> 2\sin ^{2}(2\theta) = \sin ^{2}(2\theta) + \cos ^{2}(2\theta)\] \[2\sin ^{2}(2\theta) = 1\] Using the results of trig identities the left becomes \[1 - \cos(\theta) = 1 ==> \cos(4\theta) = 0 ==> 4\theta = \frac{ \pi }{ 2 } + k \pi\]

OpenStudy (anonymous):

The last thing on the bottom left should read 1 - cos (4 theta) instead of 1 - cos (theta)